# Fast LISA likelihood approximations by downsampling

### v1

@@ 5.1 Posterior estimate evaluation times
The average PE convergence time for each of the $\Delta f/f_{\mathrm{max}}=0.9$ systems is plotted in Figure 5.1. As to be expected, the evaluation time drops as the number of datapoints is dropped. However, the times reach a minimum of a few hours at around $N_{\mathrm{s}}\approx 256$, then start to rise as $N_{\mathrm{s}}$ continues to drop. The reason for this increase is because the posterior structure begins to qualitatively break down, as extra modes are introduced and the general posterior structure becomes highly irregular, since the low number of datapoints cannot accurately encode the original posterior. The sampler therefore has significantly more difficulty in finding the isoprobability contours, and thus requires more time to converge, despite only computing a small number of datapoint values to evaluate the likelihood.
@@
Increasing $N_{\mathrm{s}}$ beyond $N_{\mathrm{s}}^{\mathrm{opt}}$ only increases the amount of time-series datapoint computations required to evaluate the likelihood, without affecting the likelihood (by definition), thus the posterior evaluation time is expected to increase roughly linearly with increasing $N_{\mathrm{s}}$. We see in Figure 5.1 that this is indeed the case, where a line of best fit is drawn through those average times for which $N_{\mathrm{s}}\geq N_{\mathrm{s}}^{\mathrm{opt}}$. Write the linear average PE time as
$T(N)$ $=aN+c$ (48a)
and for $N_{\mathrm{s}}\ll N_{\mathrm{f}}$,
$T(N_{\mathrm{s}})$ $\equiv b(2\mathcal{M}+1)N_{\mathrm{s}}+c\approx bN_{\mathcal{M}}(N_{\mathrm{s}})+c\,,$ (48b)
using (12), where we have that
$b$ $=a(2\mathcal{M}+1)^{-1}\,,$ (48c)
@@
where $T$ is the average evaluation time, $a$ is the line’s gradient and $c$ is a constant time offset that is attributed to the ‘overheads’ of the sampler (internal processes such as initialisation, computing the Fisher matrix, normalising flow transformations carried out by Nessai, etc.).
@@
The linear model of posterior evaluation time can be used to estimate the time it would take to evaluate the posterior using all the samples (when $N_{\mathcal{M}}=N_{\mathrm{Nyq}}$) from which we can roughly gauge the amount of time saved by employing our downsampling procedure. Simulated analyses of LISA data are often done in the frequency domain because of the great simplicity of computing the inner product that is afforded by the independence of the data samples there. The comparison of evaluation times shall not be especially precise because of this (the number of numerical operations required to compute signal templates will differ slightly between time domain and frequency domain signal models). Furthermore, the time domain analysis we have undertaken here requires a few more simple mathematical operations (see Section 3.1.3) in the form of element-wise products. It is reasonable to assume the cost of these operations are quite negligible in relation to the cost of evaluating the signal at the required points however, and particularly if one were to use more realistic GW signal models defined more precisely, consisting of many more terms in the expression of the strain model, $h(t)$. With these caveats mentioned, we can see the linear PE evaluation time parameters in Table 3, along with the comparison between the expected evaluation time of the original posterior with the downsampled posterior. Note that in Table 3 we have
@@
Table 3: Details of the linear best fit model of posterior evaluation times. The parameters $a,b,c$, are from Equations (5.1). Also shown is the estimated original posterior evaluation time minus the overhead time, $c$, (hours), $\Delta T_{N_{\mathrm{f}}}=\Delta T(N=N_{\mathrm{f}})$ as defined in Equation (49), the average downsampled posterior (using $N_{\mathrm{s}}^{\mathrm{opt}}\!=\!362$) evaluation time minus the overhead time (hours), $\Delta T_{N_{\mathrm{s}}^{\mathrm{opt}}}=\Delta T(N=362)$ and finally the ratio of these times (this is the approximate time saving factor of the posterior evaluation time, after disregarding the constant overheads). Recall these values are computed by setting $\mathcal{M}=7$ for all systems, the highest $\mathcal{M}$ of all systems. Most systems had $\mathcal{M}=4$, however, where $\Delta T_{N_{\mathrm{s}}}$ would be $\tfrac{3}{5}$ of the value given here, had we not manually set $\mathcal{M}=7$.
@@ 5.2 Discussion
As we can see from the rightmost column of Table 3, the reduction in evaluation times is very large. These times are computed using $N_{\mathrm{s}}=362$ and $\mathcal{M}=7$. For some systems, convergence was achieved in as little as 16 samples, and with true MCS $\mathcal{M}=4$. Recalculating by using these values for the $N_{\mathrm{f}}=10^{8}$ systems we find that $\Delta T_{N_{\mathrm{f}}}/\Delta T_{N_{\mathrm{s}}^{\mathrm{opt}}}=1841.62\to 694444.44$; i.e., likelihood evaluations in this extreme case are computed almost $7\times 10^{5}$ times faster than those of the original, fully sampled datasets would be (this is disregarding the initialisation and sampler overheads). However, most LISA signals will not have as many Nyquist datapoints, and many signals will have $\mathcal{M}>4$. Furthermore, since it is not yet known prior to study of a given PE problem what the optimal downsampling rate for that system is, it is necessary to use a higher number of datapoints to avoid potential misrepresentation of the true posterior; see Section 6.2 for further discussion on proper use of dolfen.
@@
The dimensionality of the signal model has not yet been discussed. The $n$-dimensional posteriors we have studied can be thought of as slices of $(n+1)$-dimensional posteriors. In an $(n+1)$-dimensional PE situation, the likelihood values at points on these $n$-dimensional slices will not be affected simply by allowing likelihood values on neighbouring slices also to be computed. Since there is nothing special about a particular slice, the neighbouring slices can thus be expected to be similarly accurately approximated, and one should be able to build up an accurate $(n+1)$-dimensional posterior using the same, downsampled subset of datapoints, by combining slices. One can continue to add more parameters in this way, and downsampling is not expected to be very strongly dependent on the number of free model parameters. Accuracy must of course be checked by the user for very high numbers of parameters. Some informal tests on 17-dimensional posteriors using $N_{\mathrm{s}}=500$ (default dolfen setting) have shown extremely consistent results, indicating high accuracy up to at least 17-dimensions.
We have suggested and tested various downsampling methods and schemes. After originally coming to the conclusion that hybrid, single factor downsampling was marginally more efficient in Section 4.6.1, we produced large sets of posteriors using this method in Section 5.1. We then discovered the surprising behaviour of some systems as highlighted in Figure 2 and Table 4.6.2, and recognised potential model dependent causes that might be remedied by using the FIM preservation method. A clue that FIM preservation is the more effective method was present in Figure 4.6.1, and was unfortunately missed: in that figure we can see general agreement between all methods and schemes, apart from the cluster scheme. This performed badly with single factor, but vastly improved with FIM preservation; given a datapoint subset that poorly represent the posterior under single factor modification, vast improvements can be made using FIM preservation instead. The FIM preservation statistics for a few test systems were computed and shown as the red marks in Figure 4.6.1, which at $N_{\mathrm{s}}=128$ are already approximately equal to the statistics of the converged posteriors under single factor noise reduction ($N_{\mathrm{s}}\gtrsim 1448$).
It has been demonstrated here that downsampling is a viable approach for cheaply and accurately reproducing the likelihood function in simulated environments. However, unfortunately no means of determining $N_{\mathrm{s}}^{\mathrm{opt}}$ for a given PE problem appears to be derivable from simpler objects, such as the FIM, from within the standard PE framework. For readers familiar with machine learning techniques, one can think of the downsampled data space as being akin to the so-called latent space that is often used as a low-dimensional space of auxiliary coordinates, which represents and stands in for high-dimensional data. The map from the latent space variables to the appropriate representation of the high-dimensional data is ‘learned’ by the machine learning algorithms. Downsampling seeks generality, and to skip the map learning part, by using restrictions of its ‘latent space’ (the downsampled data space) map provided by the signal model, and optimisations of that map by the downsampling methods detailed in this paper. We also mentioned that a helpful way to think of downsampling is by considering the signal and noise model at the remaining data points as an encoding scheme for the morphology of the posterior. Without first knowing the posterior however, we do not know what to encode. Downsampling is fundamentally limited by these problems, and the accuracy of a posterior can never be guaranteed, however, using the FIM preservation method over minimising Jeffreys’ divergences of the PCMs has provided a much improved and highly accurate result for our example systems. For many applications, the default settings of the downsampling software will likely yield accurate results, and should most certainly be acceptable for making informal research enquiries. However, in the absence of a guarantee of accuracy, when a user wishes to show certainty that a downsampled posterior is very highly accurate (say, for publication), it should be ensured that convergence has been reached by producing a number of posteriors at different $N_{\mathrm{s}}$ and comparing these, to show that the posteriors are independent of $N_{\mathrm{s}}$.
@@
Whilst we have seen vast improvement in PE evaluation times, many realistic signal models will be more intricate than our standard BBH inspiral signal, with likelihood functions that are generally more costly to compute. The approximation techniques like the one expounded here, and others, such as ROQ, become essential tools for experimentation in the PE environment with LISA-like signals that are prohibitively expensive to compute. As shown in the $c$ $\Delta T_{N_{\mathrm{s}}^{\mathrm{opt}}}$ columns of Table 3, the PE evaluation time for our test systems can be mostly attributed to initialisation and the internal workings of the sampler. But it is easy to envisage that, for very expensive models, which may also require using higher $N_{\mathrm{s}}$ depending on their slow evolution status, the likelihood evaluation time may again become the primary factor. Amongst possible future steps is the investigation of whether and how the different mentioned approximation techniques can be combined, to produce an even more rapid feedback arena for studying BHB signals and their various waveform modifications in LISA.
@@ 6 The dolfen package
The downsampling likelihood function estimation package: dolfen, performs all variants of the downsampling procedure that have been derived and discussed in this article. dolfen is a pip installable python package that can easily be used in conjunction with bilby [27] as per the examples provided on dolfen’s documentation pages (https://dolfen.readthedocs.io) and github pages (https://github.com/jethrolinley/dolfen).
@@ 6.1 Future developments
Recall from Section 3.4 that numerical issues often arise in exact FIM preservation. Approximate FIM preservation is far easier to solve, and dolfen appears to succeed most often in implementing this technique. In a subsequent version of dolfen, we intend to include another FIM preservation method; rather than find an individual sample reweighting solution for (approximate) FIM preservation in a given basis (see equation (26)), we shall simply use some (pre-existing) optimisation method to find individual datapoint reweightings (strictly positive) that minimises divergence between the FIMs.
@@ 6.2 Recommended usage
In order to conclusively state that a posterior approximation is close to the original posterior one is trying to estimate, the original posterior must already be known so they can be compared. Of course if one has the original, then one doesn’t need an approximation. Downsampling makes an educated guess at how best to optimise the encoding scheme provided by the random sample selection, which then can be thought of as giving a lossy compression of the likelihood. Therefore (although accuracy is likely) there are no guarantees that the likelihood returned will be accurate; the user must be responsible for this. In dolfen, one can set $N_{\mathrm{s}}$, and override the computed MCS, $\mathcal{M}$. So to begin, set, say, $N_{\mathrm{s}}=100$, and $\mathcal{M}=0$. The likelihood should not be expected to be accurate, but one can quickly get a good idea of the general likelihood morphology, to set priors and sampler settings, and so on. Once one has a rough idea of the posterior from these settings, we can use dolfen with its default settings to obtain a likely accurate likelihood. Produce a few, say 3, posteriors with different random datapoint selections (with a zero noise realisation) to confirm convergence. It is very unlikely that multiple poor representations will yield the same or similar posterior by chance, so if the posteriors are similar, convergence is clearly demonstrated. Otherwise, the number of datapoints should be increased.

### v2

@@ 5.1 Posterior estimate evaluation times
The average PE convergence time for each of the $\Delta f/f_{\mathrm{max}}=0.9$ systems is plotted in Figure 5.1. As to be expected, the evaluation time drops as the number of data points decreases. However, convergence time reaches a minimum of a few hours at around $N_{\mathrm{s}}\approx 256$, and begins to rise as $N_{\mathrm{s}}$ continues to decrease. This is because the posterior structure begins to qualitatively break down, as extra modes are introduced and the general posterior structure becomes highly irregular, since the low number of data points cannot accurately encode the original posterior. The sampler therefore has significantly more difficulty in finding the isoprobability contours, and thus requires more time to converge, despite only computing a small number of data point values to evaluate the likelihood.
@@
Increasing $N_{\mathrm{s}}$ beyond $N_{\mathrm{s}}^{\mathrm{opt}}$ only increases the amount of time-series data point computations required to evaluate the likelihood, without affecting the likelihood (by definition), thus the posterior evaluation time is expected to increase roughly linearly with increasing $N_{\mathrm{s}}$. We see in Fig. 5.1 that this is indeed the case; a line of best fit is drawn through those average times for which $N_{\mathrm{s}}\geq N_{\mathrm{s}}^{\mathrm{opt}}$. We model the mean PE time as linear in the number of (effective) samples:
T(N)=aN+c,T(N)=aN+c\,, (37)
@@
where $T$ is the average evaluation time, $N$ is the number of retained time–domain samples, $a$ is the slope and $c$ is a constant time offset (attributed to sampler overheads: initialisation, Fisher–matrix computation, normalising–flow training in Nessai, etc.).
Using Eq. (12), the number of residual points that must be computed for a given downsampling choice is $N_{\mathcal{M}}(N_{\mathrm{s}})\approx(2\mathcal{M}+1)N_{\mathrm{s}}$. Rewriting $T$ as a function of $N_{\mathcal{M}}$ gives
$T(N_{\mathrm{s}})$ $\approx aN_{\mathrm{s}}+c=a\frac{N_{\mathcal{M}}(N_{\mathrm{s}})}{2\mathcal{M}+1}+c\equiv b\,N_{\mathcal{M}}(N_{\mathrm{s}})+c\,$, (38)
so that $b=a(2\mathcal{M}+1)^{-1}$. Here $b$ is the effective slope when time is expressed as a function of the number of computed residuals $N_{\mathcal{M}}$ rather than the retained Nyquist samples $N_{\mathrm{s}}$.
@@
The linear model of posterior evaluation time can be used to estimate the cost of evaluating the likelihood on all samples ($N_{\mathcal{M}}=N_{\mathrm{Nyq}}$), and hence to gauge the time saved by downsampling. Simulated LISA analyses are often performed in the frequency domain, where the inner product is especially cheap because the data samples are independent. A direct comparison of time-domain and frequency-domain evaluation times is therefore only approximate: the number of numerical operations required to compute templates differs slightly between the two, and our time-domain scheme includes a few additional element-wise products (Sec. 3.1.3). However, these extra operations are negligible compared to the cost of evaluating realistic waveforms $h(t)$, particularly for more detailed signal models with many terms. With these caveats in mind, Table 3 reports the fitted linear parameters for the time-domain evaluation cost and compares the expected evaluation time of the original posterior with that of the downsampled posterior.
Note that in Table 3 we have
@@
Table 3: Details of the linear best fit model of posterior evaluation times. The parameters $a,b,c$, are from Equations (37) & (38). Also shown is the estimated original posterior evaluation time minus the overhead time, $c$, (hours), $\Delta T_{N_{\mathrm{f}}}=\Delta T(N=N_{\mathrm{f}})$ as defined in Equation (39), the average downsampled posterior (using $N_{\mathrm{s}}^{\mathrm{opt}}\!=\!362$) evaluation time minus the overhead time (hours), $\Delta T_{N_{\mathrm{s}}^{\mathrm{opt}}}=\Delta T(N=362)$ and finally the ratio of these times (this is the approximate time saving factor of the posterior evaluation time, after disregarding the constant overheads). Recall these values are computed by setting $\mathcal{M}=7$ for all systems, the highest $\mathcal{M}$ of all systems. Most systems had $\mathcal{M}=4$, however, where $\Delta T_{N_{\mathrm{s}}}$ would be $\tfrac{3}{5}$ of the value given here, had we not manually set $\mathcal{M}=7$.
@@ 5.2 Discussion
The rightmost column of Table 3 shows that the reduction in convergence time can already be substantial for the test systems considered here. The quoted times correspond to data lengths down to $N_{\mathrm{f}}\sim 10^{6}$, which is conservative compared to the longest LISA inspirals in a 10-year mission, which can reach $N_{\mathrm{f}}\sim 10^{9}$. The tabulated speed-ups are computed using $N_{\mathrm{s}}=362$ and $\mathcal{M}=7$. For some systems, however, convergence was achieved with as few as $16$ samples and a true MCS of $\mathcal{M}=4$. Extrapolate to $N_{\mathrm{f}}=10^{9}$ using these values, we obtain $\Delta T_{N_{\mathrm{f}}}/\Delta T_{N_{\mathrm{s}}^{\mathrm{opt}}}\approx 7\times 10^{6}$, i.e. likelihood evaluations could, in such cases, be up to $\sim 7\times 10^{6}$ times faster than for the original datasets.
In practice, most LISA signals will have fewer Nyquist data points and many will satisfy $\mathcal{M}>4$. Moreover, since the optimal downsampling rate is not known a priori, one typically adopts a somewhat larger $N_{\mathrm{s}}$ to avoid misrepresenting the true posterior (recommended usage of Dolfen in this regard is discussed in Sec. 6.2). Even so, for a 4-year LISA mission, speed-ups of order $10^{4}$ and above are often realistic.
@@
We have so far focused on $n$-dimensional posteriors that can be regarded as slices of an $(n+1)$-dimensional posterior. In an $(n+1)$-dimensional PE problem, the likelihood values on a given $n$-dimensional slice are unchanged by also evaluating the likelihood on neighbouring slices. Since no slice is special, one can expect neighbouring slices to be approximated with similar accuracy, allowing an accurate higher-dimensional posterior to be built up from the same downsampled subset of data points. In this sense, downsampling is not expected to depend strongly on the number of free model parameters, although accuracy should always be checked for very high dimensionality. Informal tests on 17-dimensional posteriors using $N_{\mathrm{s}}=500$ (the default Dolfen setting) have shown extremely consistent results, indicating excellent accuracy up to at least 17 parameters.
We have proposed and tested several downsampling methods and schemes. In Sec. 4.6.1 we initially found that hybrid single-factor downsampling was marginally more efficient, and we therefore used this method to produce large sets of posteriors in Sec. 5.1. Subsequent investigation of the surprising behaviour seen in Fig. 2 and Table 4.6.2 led us to identify model-dependent effects that can be mitigated by using the FIM-preservation method. In hindsight, Fig. 4.6.1 already contained a clue: all methods and schemes broadly agree, except for the cluster scheme, which performs poorly with single-factor downsampling but improves dramatically when FIM preservation is used. The red markers in Fig. 4.6.1 show that, at $N_{\mathrm{s}}=128$, the FIM-preservation statistics are already comparable to those of the converged posteriors obtained with single-factor noise reduction ($N_{\mathrm{s}}\gtrsim 1448$).
Our results demonstrate that downsampling is a viable way to reproduce the likelihood function cheaply and accurately in simulated environments. However, we have not found any reliable method of predicting $N_{\mathrm{s}}^{\mathrm{opt}}$ for a given PE problem from simpler diagnostic quantities (such as the FIM) within the standard framework. For readers familiar with machine learning, it is useful to view the downsampled data space as analogous to a latent space: a reduced representation that stands in for the full high-dimensional data. In machine learning, the map from latent variables to data is learned; here, downsampling instead exploits restrictions of a map already provided by the signal model, refined by the optimisation procedures described in this paper. Equivalently, one may regard the signal and noise model evaluated at the retained data points as an encoding of the posterior geometry. Without prior knowledge of the posterior, however, we do not know in advance what should be encoded, and downsampling is fundamentally limited by this; perfect accuracy can never be guaranteed. Nevertheless, using FIM preservation rather than minimising Jeffreys divergences of the PCMs yields substantially improved and highly accurate results for our example systems.
For most applications, the default Dolfen settings will likely produce sufficiently accurate posteriors and should be more than adequate for exploratory or informal investigations. When a high level of certainty is required, we recommend explicitly checking convergence by generating posteriors at several values of $N_{\mathrm{s}}$ and verifying that the resulting posteriors are effectively independent of $N_{\mathrm{s}}$ within the admissible range.
@@
Although we have seen dramatic improvements in PE evaluation times, many realistic signal models will be more intricate than our standard BHB inspiral signal, with likelihood functions that are more expensive to compute. Approximation techniques such as the one presented here, and others like ROQ, will be essential tools for experimentation with LISA-like signals whose full likelihoods are prohibitively costly to evaluate. A natural direction for future work is to investigate how these approaches can be combined to provide an even more rapid feedback arena for studying BHB signals and their waveform modifications in the LISA context.
Finally, in applications that aim to quantify systematic modelling bias, one may precompute the Fisher information matrix for the injected signal (generated with the “true” physics) and use it to define the downsampled inner product. In this setup, the downsampled data preserve the (Fisher) information content of the original signal about the true parameters, while the same inner product is then used to evaluate the likelihood under a mismatched template model. This is a natural strategy for bias studies, but the resulting likelihood still needs to be validated for the chosen template. The key advantage is that the approximation scheme is held fixed, so any observed bias in parameter recovery can be attributed to the physics mismatch itself, rather than to changes in the numerical representation of the data.
@@ 6 The Dolfen package
The downsampling likelihood function estimation package Dolfen, performs all variants of the downsampling procedure that have been derived and discussed in this article. Dolfen is a pip installable Python package that can easily be used in conjunction with Bilby [Bilby] as per the examples provided on Dolfen’s documentation pages (https://dolfen.readthedocs.io) and github pages (https://github.com/jethrolinley/dolfen).
@@ 6.1 Future developments
Recall from Section 3.4 that numerical issues often arise in exact FIM preservation. Approximate FIM preservation is far easier to solve, and Dolfen appears to succeed most often in implementing this technique. In a subsequent version of Dolfen, we intend to include another FIM preservation method; rather than find an individual sample reweighting solution for (approximate) FIM preservation in a given basis (see equation (66)), we shall use a (pre-existing) numerical optimisation method to find individual data point reweightings (strictly positive) that minimises divergence between the FIMs.
Real LISA streams will contain short veto windows (telemetry drop-outs, thruster events, etc.). In the present release a user can already approximate such data by (i) cutting the timeline into $M$ contiguous, gap-free segments, (ii) running dolfen.log_likelihood on each segment independently, and (iii) summing the log-likelihoods. Because the segments are disjoint, this is exactly the “cluster-sampling” factorisation introduced in Section 3.1.1, with clusters chosen by the user rather than by the current fixed method. The next version of Dolfen will automate this procedure, where the user can list time intervals to be excluded.
@@ 6.2 Recommended usage
To assert that an approximate posterior is close to the true posterior, one would in principle need access to the true posterior itself in order to compare them directly. Of course, if the true posterior were already known, no approximation would be required. Downsampling instead makes an informed attempt to optimise the encoding of information contained in a randomly selected subset of data points, yielding a lossy compression of the likelihood. As with any likelihood approximation, its accuracy must be assessed in context.
We therefore recommend treating downsampling as a controlled approximation and checking convergence with respect to the number of retained samples, $N_{\mathrm{s}}$, and the resulting MCS, $\mathcal{M}$. In Dolfen, both $N_{\mathrm{s}}$ and $\mathcal{M}$ can be explicitly set, overriding the automatically computed MCS if desired. A practical workflow is:
  1. 1.
    Exploratory run. Start with a deliberately small choice, e.g. $N_{\mathrm{s}}=100$ and $\mathcal{M}=1$. The resulting likelihood should not be expected to be highly accurate, but it is typically sufficient to explore the rough likelihood geometry, tune priors, and adjust sampler settings.
  2. 2.
    Default (high-accuracy) runs. Once a sensible configuration has been identified, run Dolfen with its default settings to obtain a likely accurate likelihood.
  3. 3.
    Convergence check. Generate several (e.g. three) posterior samples using different random selections of data points (with zero noise realisations) and different values of $N_{\mathrm{s}}$ (for example, $N_{\mathrm{s}}=300,600,900$). If the resulting posteriors are mutually consistent, this provides strong evidence that the approximation has converged. If not, increase $N_{\mathrm{s}}$ and repeat.
In practice, it is highly implausible that several genuinely poor and mutually distinct likelihood approximations would, by chance, yield closely similar posteriors. Consequently, stability of the posterior under changes in $N_{\mathrm{s}}$ and random subsampling provides a practical and robust diagnostic of approximation accuracy.